# Glassy dynamics in deep neural networks: A structural comparison

### v2

@@ Appendix B Choosing $\epsilon$
Appendix C Measurement process
In Fig. A2 we demonstrate a typical training loss curve illustrating the process of taking measurements. An initial period of SGD moves the network out of its initial condition and into a local minimum. The dynamics then switch to Eq. (3) (Langevin), there is an initial transient response to the new dynamics, and after some time $t_{w}$, taken to be 20000 timesteps unless otherwise stated, measurement starts. The choice of $t_{w}=20000$ was made somewhat subjectively, balancing the fact that a longer $t_{w}$ reaches a flatter loss curve, but we only have finite training time.
Figure A2: The training loss curve of a network, illustrating the measurement process.
@@ Appendix I Non-Gaussianity and Dynamic Heterogeneity in CIFAR-10
Using the DNNs trained on CIFAR-10 from Sec. E, we measure the non-Gaussian parameter, $\alpha$, and the distribution of weight motilities at a range of temperatures. These networks exhibit the same behaviour as the MNIST examples in the main text. All three phases have $\alpha\neq 0$, starting at long times for high temperatures, then occurring at all times for low temperatures. This behaviour is less pronounced for the overparameterised phase. Note that the pronounced oscillations up to $t=100$ are due to the fact that the network states are saved at a higher frequency at shorter times. The weight motilities follow bimodal distributions, i.e. demonstrate DH, at low temperatures for the regularised and underparameterised, but not overparameterised, phases.
Figure A8: The non-Gaussian parameter, $\alpha$, and weight motility distributions for DNNs trained on CIFAR-10. Non-Gaussian behaviour, corresponding to $\alpha\neq 0$ occurs at long times for high temperatures, and all times for low temperatures in the regularised (a), underparameterised (b) and overparameterised (c) phases. The temperatures in (a) and (b), going from blue (high curves) to red (low curves) are $T=10^{-10}$, $10^{-8}$, $10^{-6}$, $10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$. The temperatures in (c) are $T=10^{-10}$, $10^{-8}$, $10^{-6}$, $10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $1$. The regularised and underparameterised networks also have bimodal weight motility distributions (panels (d) and (e) respectively), unlike the overparameterised phase (f). The temperatures used in (d), (e), and (f), going from blue (left curves) to red (right curves) are $T=10^{-8}$, $10^{-6}$, $10^{-5}$, $10^{-3}$.

### v3

@@ Appendix I Non-Gaussianity and Dynamic Heterogeneity in CIFAR-10
Using the DNNs trained on CIFAR-10 from Sec. E, we measure the non-Gaussian parameter, $\alpha_{2}$, and the distribution of weight motilities at a range of temperatures. These networks exhibit the same behaviour as the MNIST examples in the main text. All three phases have $\alpha_{2}\neq 0$, starting at long times for high temperatures, then occurring at all times for low temperatures. This behaviour is less pronounced for the overparameterised phase. Note that the pronounced oscillations up to $t=100$ are due to the fact that the network states are saved at a higher frequency at shorter times. The weight motilities follow bimodal distributions, i.e. demonstrate DH, at low temperatures for the regularised and underparameterised, but not overparameterised, phases.
Figure A8: The non-Gaussian parameter, $\alpha_{2}$, and weight motility distributions for DNNs trained on CIFAR-10. Non-Gaussian behaviour, corresponding to $\alpha_{2}\neq 0$ occurs at long times for high temperatures, and all times for low temperatures in the regularised (a), underparameterised (b) and overparameterised (c) phases. The temperatures in (a) and (b), going from blue (high curves) to red (low curves) are $T=10^{-10}$, $10^{-8}$, $10^{-6}$, $10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $10^{-1}$, $1$. The temperatures in (c) are $T=10^{-10}$, $10^{-8}$, $10^{-6}$, $10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $1$. The regularised and underparameterised networks also have bimodal weight motility distributions (panels (d) and (e) respectively), unlike the overparameterised phase (f). The temperatures used in (d), (e), and (f), going from blue (left curves) to red (right curves) are $T=10^{-8}$, $10^{-6}$, $10^{-5}$, $10^{-3}$.